# Adaptive Rollout Allocation for Online Reinforcement Learning with Verifiable Rewards

### v1

@@ Abstract
Sampling efficiency is a key bottleneck in reinforcement learning with verifiable rewards. Existing group-based policy optimization methods, such as GRPO, allocate a fixed number of rollouts for all training prompts. This uniform allocation implicitly treats all prompts as equally informative, and could lead to inefficient computational budget usage and impede training progress. We introduce VIP, a Variance-Informed Predictive allocation strategy that allocates a given rollout budget to the prompts in the incumbent batch to minimize the expected gradient variance of the policy update. At each iteration, VIP uses a lightweight Gaussian process model to predict per-prompt success probabilities based on recent rollouts. These probability predictions are translated into variance estimates, which are then fed into a convex optimization problem to determine the optimal rollout allocations under a hard compute budget constraint. Empirical results show that VIP consistently improves sampling efficiency and achieves higher performance than uniform or heuristic allocation strategies in multiple benchmarks. Our code will be available at https://github.com/HieuNT91/VIP
@@ 1 Introduction
The advent of Language Models (LMs) has marked a new era in artificial intelligence, transforming models from static knowledge bases into dynamic, intelligent agents capable of complex reasoning, creative generation, and intricate problem-solving. This evolution has been largely driven by the advancement of post-training methodologies, which are crucial for aligning a base LM’s vast, general knowledge with specific human intentions, ethical guidelines, and desired task-oriented behaviors. These refinement processes bridge the gap between a model that understands language and one that can reliably and safely act as a collaborator, assistant, or autonomous agent. As LMs become more integrated into real-world applications, the effectiveness and efficiency of these post-training stages have become a primary focus of research and development.
Two prominent paradigms are used for post-training LMs: Supervised Fine-Tuning (SFT) and Reinforcement Learning (RL). SFT is a computationally efficient approach that trains the model on a curated dataset of input-output pairs, effectively equipping it with domain-specific knowledge. However, its effectiveness is often hampered by the difficulty of curating high-quality and comprehensive datasets. In contrast, RL-based methods, which treat the model as an agent optimizing for a reward signal derived from human preferences (Reinforcement Learning from Human Feedback, RLHF) or external verification mechanisms (Reinforcement Learning from Verifiable Rewards, RLVR), can explore a broader action space and achieve superior performance on complex, open-ended tasks. RLVR, in particular, has gained wide popularity thanks to its low reliance on human input. However, the performance of RL methods comes at a significant computational cost. Algorithms like Proximal Policy Optimization (PPO, Schulman et al. (2015)) require the simultaneous training of a separate value model to estimate the advantage of actions, which can be prohibitively memory-intensive in resource-constrained situations. To address this, a family of “critic-free” or group-based RL methods has emerged, estimating the advantage as a relative measure within a group of rollouts, essentially trading speed for memory efficiency. Such methods include GRPO, (Shao et al., 2024), Dr. GRPO (Liu et al., 2025), RLOO (Ahmadian et al., 2024), and other variants. A direct drawback of these methods is the additional computation time required to generate multiple rollouts for a training example. The number of generations often needs to be large enough (e.g., 16) to obtain a stable training process, which exacerbates the generation overhead and eventually leads to a memory-bound system. Such high computational demands consequently highlight the need for greater sampling efficiency in the RL training process.
@@ 2 Related Work
Recent advancements in group-based RL for LMs have focused on adaptive strategies to enhance both training efficiency and performance. These methods move beyond static training pipelines by dynamically selecting and managing data used for policy updates. Initial research empirically demonstrated that filtering out “non-informative” problems to which the model’s accuracy is either 0 or 1 can improve training efficiency, as these problems lead to rollout batches with zero variance and consequently make no contribution to the gradient signal (Yu et al., 2025). Along similar lines, (Lin et al., 2025) filtered out rollouts with low absolute advantage, while (Xu et al., 2025) proposed to retain examples with the highest variance. Lacking an a priori measure of training prompt informativeness, these methods require the pre-sampling of an large batch of rollouts before each update iteration, potentially negating efficiency gains from the significant sampling overhead.
@@ 3 Background on RLVR
For prompt $q$, the term $A^{\prime}_{j,\tau}$ is the normalized or advantage-shaped reward at token $\tau$ for output $\tilde{o}_{j}$:
@@ 4 Analysis for Gradient Variance
Above, $H(\tilde{o}_{j})$ is the average log-likelihood gradient, where the average is taken over all tokens in the $j$-th rollout. Consider a specific prompt $q$ and the $n$ rollouts $\{\tilde{o}_{k}\}$, the contribution of $q$ to the calculation of the sample gradient is $\frac{1}{n}\sum_{j=1}^{n}A_{j}(\{\tilde{o}_{k}\})H(\tilde{o}_{j})$, and we omit the subscript $q$ to avoid clutter. This quantity is a random vector because $H(\tilde{o}_{j})$ is a random vector. We will study the projected gradient, obtained by projecting $H(\tilde{o}_{j})$ onto the vector $\mathbbm{1}$ of all ones. To this end, we are interested in the (univariate) random variable:
$\tilde{G}\triangleq\frac{1}{n}\sum\nolimits_{j=1}^{n}A_{j}(\{\tilde{o}_{k}\})\mathbbm{1}^{\top}H(\tilde{o}_{j}).$
To simplify the notation, we define the random variables $\tilde{R}_{j}=R(\tilde{o}_{j})$ and $\tilde{Z}_{j}=\mathbbm{1}^{\top}H(\tilde{o}_{j})$ for all $j$. We first consider the Dr. GRPO case. The gradient $\tilde{G}$ has the specific form:
@@
Assumption 4.1 is realistic in common settings: given the same prompt and fixed sampling procedure, different rollouts are independent samples from the same distribution $\pi_{\theta}$. Thus, it is reasonable to assume that $\{\tilde{o}_{j}\}$ are i.i.d., which induces Assumption 4.1(i). Assumption 4.1(ii) needs further justification because for the $j$-th rollout, the reward $\tilde{R}_{j}$ may be correlated with the average gradient $\tilde{Z}_{j}$. However, we have strong reasons to believe that it holds because $\tilde{R}$ is a Bernoulli distribution, and $\tilde{Z}_{j}$ has a doubly-averaging nature over the number of tokens (from the definition of $H(\tilde{o}_{j})$) and over the number of dimensions in the gradient vector (from $\mathbbm{1}^{\top}H(\tilde{o}_{j})$). Empirically, we include in Appendix B a statistical test to support our assumption.
@@ 6 Numerical Experiments
We now showcase the empirical benefit of our VIP framework on the mathematical reasoning task and the tool-augmented reasoning task. The common setup for both tasks is as follows: we encode each prompt $q$ into a 384-dimensional vector $x_{q}$ using all-MiniLM-L6-v2. Before training, we pre-compute and cache the pairwise Euclidean distances between all prompt pairs, making the runtime overhead negligible. We use a median heuristic to set the bandwidth $h$ of the kernel $\mathcal{K}$. We also assume that $\tilde{Z}_{q}$ has the same variances over all $q$ when we compute the allocation; we support this assumption with a statistical test presented in Appendix B.3. We train all models for one epoch under two total rollout budgets, $6\times Q$ and $8\times Q$, where $Q$ is the dataset size. We integrate VIP on top of two group-relative policy optimization baselinesm, RLOO and Dr. GRPO, implemented using VERL (Sheng et al., 2024) for math experiments and following  (Chen et al., 2025b; Jin et al., 2025) for tool-augmented tasks. We closely follow the default hyperparameter settings used in RL training. The anonymized repository is https://anonymous.4open.science/r/VIP-2F00.
@@
Qwen2.5-Math-1.5B Dr. GRPO 22.0 4.0 7.0 25.0 3.0 8.0
@@ 7 Conclusion
Reproducibility Statement. We have made every effort to ensure that our results are reproducible. All of the models, algorithms, and experiments described in the paper are accompanied by detailed explanations in the main text and appendix. To further support reproducibility, we have provided an anonymized repository containing the full source code, along with instructions for running the experiments and replicating our results (see the link in the supplementary materials). The datasets we used are publicly available, and we include clear descriptions of any preprocessing steps in the appendix. For all theoretical results, we have stated our assumptions explicitly and included complete proofs in the supplementary material.
@@
Acknowledgments. Viet Anh Nguyen gratefully acknowledges the support from the CUHK’s Improvement on Competitiveness in Hiring New Faculties Funding Scheme, UGC ECS Grant 24210924, and UGC GRF Grant 14208625.
@@ A.1 Proofs of Section 4
$=\frac{1}{n^{2}}\mathbb{E}\left[\sum_{j=1}^{n}\tilde{R}_{j}^{2}\tilde{Z}_{j}^{2}+\sum_{1\leq j<k\leq n}\tilde{R}_{j}\tilde{Z}_{j}\tilde{R}_{k}\tilde{Z}_{k}\right]$
$=\frac{1}{n^{2}}\left(\sum_{j=1}^{n}\mathbb{E}\left[\tilde{R}_{j}^{2}\tilde{Z}_{j}^{2}\right]+\sum_{1\leq j<k\leq n}\mathbb{E}\left[\tilde{R}_{j}\tilde{Z}_{j}\tilde{R}_{k}\tilde{Z}_{k}\right]\right)$
$=\frac{1}{n^{2}}\left(\sum_{j=1}^{n}\mathbb{E}\left[\tilde{R}_{j}^{2}\tilde{Z}_{j}^{2}\right]+\sum_{1\leq j<k\leq n}\mathbb{E}\left[\tilde{R}_{j}\tilde{Z}_{j}\right]\mathbb{E}\left[\tilde{R}_{k}\tilde{Z}_{k}\right]\right)$ $(\text{independence between $j$ and $k$})$
@@ Appendix F Training Evolution Comparison
In this section, we assess the robustness and stability of our method by retraining Qwen2.5-Math-1.5B using GRPO, RLOO, and their VIP-augmented counterparts (GRPO+VIP, RLOO+VIP) across five random seeds. Figures 6 and 7 report the mean and standard deviation for multiple performance metrics (best@32, maj@32, mean@32).
To ensure that all training trajectories are directly comparable, every model is trained on the same dataset under identical optimization settings: the same fixed ordering of 17k training prompts, one epoch of training, a batch size of 512, mini-batch size of 64, and rollout budget per batch of 512 * 8. As a result, each gradient step corresponds to the same amount of data and computation across all methods.
Across all seeds and evaluation checkpoints, we observe consistent and pronounced improvements from using VIP:
[…]
(ii) Steeper and more reliable improvement per gradient step. VIP consistently increases the slope of the learning curve. Its trajectories rise smoothly and monotonically, while the baselines (particularly GRPO on AIME2025 best@32) often progress slowly or temporarily plateau between steps 10–20. This shows that variance-aware allocation accelerates the effective learning rate without introducing instability.
(iii) Increased training stability. VIP reduces variance across seeds and produces smoother learning curves, reflecting more stable gradient updates. This aligns with the goal of variance-informed allocation: reducing gradient noise directly translates into more predictable and reliable optimization dynamics.
Together, these results demonstrate that VIP improves both the speed and the stability of GRPO and RLOO training, leading to faster convergence and consistently higher performance throughout the entire training trajectory.
@@
(d) AIME 2025, best@32
(e) AIME 2025, maj@32
(f) AIME 2025, mean
Figure 6: GRPO vs. GRPO+VIP on AIME 2024 and 2025 across different accuracy metrics.
(a) AIME 2024, best@32
(b) AIME 2024, maj@32
(c) AIME 2024, mean
(d) AIME 2025, best@32
(e) AIME 2025, maj@32
(f) AIME 2025, mean
Figure 7: RLOO vs. RLOO+VIP on AIME 2024 and 2025 across different accuracy metrics.

### v3

@@ Abstract
Sampling efficiency is a key bottleneck in reinforcement learning with verifiable rewards. Existing group-based policy optimization methods, such as GRPO, allocate a fixed number of rollouts for all training prompts. This uniform allocation implicitly treats all prompts as equally informative, and could lead to inefficient computational budget usage and impede training progress. We introduce VIP, a Variance-Informed Predictive allocation strategy that allocates a given rollout budget to the prompts in the incumbent batch to minimize the expected gradient variance of the policy update. At each iteration, VIP uses a lightweight Gaussian process model to predict per-prompt success probabilities based on recent rollouts. These probability predictions are translated into variance estimates, which are then fed into a convex optimization problem to determine the optimal rollout allocations under a hard compute budget constraint. Empirical results show that VIP consistently improves sampling efficiency and achieves higher performance than uniform or heuristic allocation strategies in multiple benchmarks.
@@ 1 Introduction
The advent of Language Models (LMs) has marked a new era in artificial intelligence, transforming models from static knowledge bases into dynamic, intelligent agents capable of complex reasoning, creative generation, and intricate problem-solving. This evolution has been largely driven by the advancement of post-training methods (Tong et al., 2024; Rafailov et al., 2023; Shao et al., 2024) and test-time scaling methods (Snell et al., 2024; Feng et al., 2023; Nguyen et al., 2025a; b), which are crucial for aligning a base LM’s vast, general knowledge with specific human intentions, ethical guidelines, and desired task-oriented behaviors. These refinement processes bridge the gap between a model that understands language and one that can reliably and safely act as a collaborator, simulator, or autonomous agent. As LMs become more integrated into real-world applications, the effectiveness and efficiency of these post-training stages have become a primary focus of research and development.
While test-time scaling methods are training-free and effective in many scenarios, post-training generally yields higher performance gains at a lower inference budget. Two prominent paradigms are used for post-training LMs: Supervised Fine-Tuning (SFT) and Reinforcement Learning (RL). SFT is a computationally efficient approach that trains the model on a curated dataset of input-output pairs, effectively equipping it with domain-specific knowledge. However, its effectiveness is often hampered by the difficulty of curating high-quality and comprehensive datasets. In contrast, RL-based methods, which treat the model as an agent optimizing for a reward signal derived from human preferences (Reinforcement Learning from Human Feedback, RLHF) or external verification mechanisms (Reinforcement Learning from Verifiable Rewards, RLVR), can explore a broader action space and achieve superior performance on complex, open-ended tasks. RLVR, in particular, has gained wide popularity thanks to its low reliance on human input. However, the performance of RL methods comes at a significant computational cost. Algorithms like Proximal Policy Optimization (PPO, Schulman et al. (2015)) require the simultaneous training of a separate value model to estimate the advantage of actions, which can be prohibitively memory-intensive in resource-constrained situations. To address this, a family of “critic-free" or group-based RL methods has emerged, estimating the advantage as a relative measure within a group of rollouts, essentially trading speed for memory efficiency. Such methods include GRPO, (Shao et al., 2024), Dr. GRPO (Liu et al., 2025), RLOO (Ahmadian et al., 2024), and other variants. A direct drawback of these methods is the additional computation time required to generate multiple rollouts for a training example. The number of generations often needs to be large enough (e.g., 16) to obtain a stable training process, which exacerbates the generation overhead and eventually leads to a memory-bound system. Such high computational demands consequently highlight the need for greater sampling efficiency in the RL training process.
@@ 2 Related Work
Recent advancements in group-based RL for LMs have focused on adaptive strategies to enhance both training efficiency and performance. These methods move beyond static training pipelines by dynamically selecting and managing data used for policy updates. Initial research empirically demonstrated that filtering out “non-informative" problems to which the model’s accuracy is either 0 or 1 can improve training efficiency, as these problems lead to rollout batches with zero variance and consequently make no contribution to the gradient signal (Yu et al., 2025). Along similar lines, (Lin et al., 2025) filtered out rollouts with low absolute advantage, while (Xu et al., 2025) proposed to retain examples with the highest variance. Lacking an a priori measure of training prompt informativeness, these methods require the pre-sampling of an large batch of rollouts before each update iteration, potentially negating efficiency gains from the significant sampling overhead.
@@ 3 Background on RLVR
For prompt $q$, the term $A^{\prime}_{j,\tau}$ is the clipped surrogate at token $\tau$ for output $\tilde{o}_{j}$:
@@ 4 Analysis for Gradient Variance
Above, $H(\tilde{o}_{j})$ is the average log-likelihood gradient, where the average is taken over all tokens in the $j$-th rollout. Consider a specific prompt $q$ and the $n$ rollouts $\{\tilde{o}_{k}\}$, the contribution of $q$ to the calculation of the sample gradient is $\frac{1}{n}\sum_{j=1}^{n}A_{j}(\{\tilde{o}_{k}\})H(\tilde{o}_{j})$, and we omit the subscript $q$ to avoid clutter. This quantity is a random vector because $H(\tilde{o}_{j})$ is a random vector. To make the variance analysis mathematically tractable while accurately proxying the scale of the parameter updates, we will study the $L_{2}$ norm of this gradient gradient contribution. To this end, we are interested in the (univariate) random variable:
$\tilde{G}\triangleq\frac{1}{n}\sum\nolimits_{j=1}^{n}A_{j}(\{\tilde{o}_{k}\})\|H(\tilde{o}_{j})\|_{2}.$
To simplify the notation, we define the random variables $\tilde{R}_{j}=R(\tilde{o}_{j})$ and $\tilde{Z}_{j}=\|H(\tilde{o}_{j})\|_{2}$ for all $j$. We first consider the Dr. GRPO case. The gradient $\tilde{G}$ has the specific form:
@@
Assumption 4.1 is realistic in common settings: given the same prompt and fixed sampling procedure, different rollouts are independent samples from the same distribution $\pi_{\theta}$. Thus, it is reasonable to assume that $\{\tilde{o}_{j}\}$ are i.i.d., which induces Assumption 4.1(i). Assumption 4.1(ii) needs further justification because for the $j$-th rollout, the reward $\tilde{R}_{j}$ may be correlated with the average gradient $\tilde{Z}_{j}$. However, we have strong reasons to believe that it holds because $\tilde{R}$ is a Bernoulli distribution, and $\tilde{Z}_{j}$ aggregates variations both over the number of tokens (from the definition of $H(\tilde{o}_{j})$) and across the high-dimensional parameter space (from $\|H(\tilde{o}_{j})\|_{2}$). Empirically, we include in Appendix B a statistical test to support our assumption.
@@ 6 Numerical Experiments
We now showcase the empirical benefit of our VIP framework on the mathematical reasoning task and the tool-augmented reasoning task. The common setup for both tasks is as follows: we encode each prompt $q$ into a 384-dimensional vector $x_{q}$ using all-MiniLM-L6-v2. Before training, we pre-compute and cache the pairwise Euclidean distances between all prompt pairs, making the runtime overhead negligible. We use a median heuristic to set the bandwidth $h$ of the kernel $\mathcal{K}$. We also assume that $\tilde{Z}_{q}$ has the same variances over all $q$ when we compute the allocation; we support this assumption with a statistical test presented in Appendix B.3. We train all models for two epochs under two total rollout budgets, $8\times Q$ and $16\times Q$, where $Q$ is the dataset size. We integrate VIP on top of two group-relative policy optimization baselines, RLOO and Dr. GRPO, implemented using VERL (Sheng et al., 2024) for math experiments and following  (Chen et al., 2025b; Jin et al., 2025) for tool-augmented tasks. We closely follow the default hyperparameter settings used in RL training.
@@
Qwen2.5-Math-1.5B Dr. GRPO 23.96 9.479 13.64 28.21 5.83 9.97
@@ 7 Conclusion
Reproducibility Statement. We have made every effort to ensure that our results are reproducible. All of the models, algorithms, and experiments described in the paper are accompanied by detailed explanations in the main text and appendix. To further support reproducibility, we have provided an anonymized repository containing the full source code, along with instructions for running the experiments and replicating our results. The datasets we used are publicly available, and we include clear descriptions of any preprocessing steps in the appendix. For all theoretical results, we have stated our assumptions explicitly and included complete proofs in the supplementary material.
@@
Acknowledgments. Viet Anh Nguyen gratefully acknowledges the support from the CUHK’s Improvement on Competitiveness in Hiring New Faculties Funding Scheme, UGC ECS Grant 24210924, and UGC GRF Grant 14208625. The authors would like the thank Dongxuan Zhu and Hongzheng Yang for their helpful discussion regarding Resource Allocation and Reinforcement Learning.
@@ A.1 Proofs of Section 4
$=\frac{1}{n^{2}}\mathbb{E}\left[\sum_{j=1}^{n}\tilde{R}_{j}^{2}\tilde{Z}_{j}^{2}+2\sum_{1\leq j<k\leq n}\tilde{R}_{j}\tilde{Z}_{j}\tilde{R}_{k}\tilde{Z}_{k}\right]$
$=\frac{1}{n^{2}}\left(\sum_{j=1}^{n}\mathbb{E}\left[\tilde{R}_{j}^{2}\tilde{Z}_{j}^{2}\right]+2\sum_{1\leq j<k\leq n}\mathbb{E}\left[\tilde{R}_{j}\tilde{Z}_{j}\tilde{R}_{k}\tilde{Z}_{k}\right]\right)$
$=\frac{1}{n^{2}}\left(\sum_{j=1}^{n}\mathbb{E}\left[\tilde{R}_{j}^{2}\tilde{Z}_{j}^{2}\right]+2\sum_{1\leq j<k\leq n}\mathbb{E}\left[\tilde{R}_{j}\tilde{Z}_{j}\right]\mathbb{E}\left[\tilde{R}_{k}\tilde{Z}_{k}\right]\right)$ $(\text{independence between $j$ and $k$})$
@@ Appendix F Training Evolution Comparison
In this section, we compare training evolution on model Qwen2.5-Math-1.5B using GRPO and GRPO+VIP. Figures 6 report training metrics (mean advantages, mean rewards) and multiple performance metrics (best@32, maj@32, mean@32).
To ensure that all training trajectories are directly comparable, every model is trained on the same dataset under identical optimization settings: the same fixed ordering of 17k training prompts, two epochs of training, a batch size of 512, mini-batch size of 64, and rollout budget per batch of 512 * 8 and 512 * 16. As a result, each gradient step corresponds to the same amount of data and computation across all methods.
Across most of evaluation checkpoints, we observe consistent and pronounced improvements from using VIP. During training, we observe that using VIP, we obtain consistently far-from-zero advantages and higher rewards. This indicates that our learning algorithm provides richer learning signal and our rollout allocation strategy is effective. Further empirical evidence are provided in Figures 7 where VIP performance exceeded GRPO at most training steps.
Figure 6: Mean advantages and mean rewards of GRPO vs. GRPO+VIP using rollout budgets of 8 (first row) and 16 (second row).
@@
(d) AIME 2024, best@32
(e) AIME 2024, maj@32
(f) AIME 2024, mean
(g) AIME 2025, best@32
(h) AIME 2025, maj@32
(i) AIME 2025, mean
(j) AIME 2025, best@32
(k) AIME 2025, maj@32
(l) AIME 2025, mean
Figure 7: GRPO vs. GRPO+VIP using rollout budget 8 (a,b,c,g,h,i) and 16 (d,e,f,j,k,l) on AIME 2024 and 2025 across different accuracy metrics.